Several complaints recently have been sent to the LA city office regarding the increasing incidence of congestion on the city’s streets. The complaints attribute the cause of these traffic tie-ups to a lack of synchronization of the traffic lights. The lights are controlled by a main computer system, and adjusting this program is costly. Therefore, the city is reluctant to change the situation unless a clear need is shown. During the past three months, the office initiated a study and collected 200 data each month at major intersections and identify the # of congestion incidents. The congestion data is shown below.

Month Congestion Incidents
1 12
2 18
3 15
  • Keep 3-decimal if not exact, do not round up. For example, 0.24364....will be kept as 0.243
  • Both formats, 0.243 or .243, are acceptable
  1. What is the sample size of the study?
  2. What is the p-value of sample 1 (month 1)?
  3. What is the p-bar value of the study?
  4. What is the upper control limit, for a 2-sigma quality standard, based on the above data?
  5. Assume the p-chart for an acceptable traffic congestion rate is [0.02, 0.09], and last month's # of congestions is 213 from a 2400 traffic-observation, what is last month's sample p-value?  Does last month's congestion situation acceptable or not, comparing to the assumed p-chart? (answer Y or N here)
Answer #1

Sample size for study is 200.

p value for first month = 12/200= 0.06

overall p value = (12+18+15) / 3*200 =0.075

UCLp = pbar + 2*[ pbar*(1-pbar)/n]

= 0.075 + 2*[ 0.075*(1-0.075)/200]^0.5 = 0.075+0.03724 =0.1122

Last month's p value = 213 /2400 =0.08875

which is less than 0.09 and hence acceptable